OpenStudy (anonymous):

Let f(x) = 2-x and g(x) = 1/x. Perform each function operation and then find the domain of the result. (f-g)(x) Show steps please, very confused

hero (hero):

(f-g)(x) = f(x) - g(x) = (2-x) - (1/x) = x(2-x)/x - 1/x = (2x - x^2)/x - 1/x = (2x - x^2 - 1)/x

OpenStudy (anonymous):

do you do the reciprocal of x/1?

hero (hero):

(-x^2 + 2x - 1)/x = -(x^2 - 2x + 1)/x = -(x-1)(x-1)/x Therefore, (f-g)(x) = -((x-1)(x-1))/x

OpenStudy (anonymous):

Do you think you can do it in the equation thing, i get confused looking at it like that

hero (hero):

\[(f-g)(x) = f(x) - g(x)\]\[f(x) - g(x) = (2-x) - (\frac{1}{x})\]

hero (hero):

\[f(x) - g(x) = \frac{x(2-x)}{x} - \frac{1}{x}\]\[f(x) - g(x) = \frac{2x - x^2}{x} - \frac{1}{x}\]\[f(x) - g(x) = \frac{2x - x^2 - 1}{x}\]

OpenStudy (anonymous):

why do you divide 2-x by x and multiply by x??

hero (hero):

\[f(x) - g(x) = \frac{-x^2 + 2x - 1}{x} = -\frac{(x-1)(x-1)}{x}\]

hero (hero):

We do that so that we get a common denominator of x to combine everything into one fraction. Remember, x/x = 1

OpenStudy (anonymous):

the the domain would be all real numbers? is there any exceptions?

hero (hero):

Maybe I should have wrote it like this so that it is easier for you to see: \[\frac{x}{x} \times \frac{2-x}{1} = \frac{x(2-x)}{x}\]

hero (hero):

\[x \ne 0\]

hero (hero):

That literally means "x cannot equal zero"

OpenStudy (anonymous):

you always multiply by x/x when the second equation is a fraction that has x?

hero (hero):

If you want to combine a number with a fraction that has denominator x, then yes.

hero (hero):

If the fraction had denominator 2x, then you would multiply the number by 2x/2x

hero (hero):

How would you add \[6 + \frac{6}{7}\]

OpenStudy (anonymous):

so something like (3-x) - (3/6x) would mean you multiply by 6x/6x?

hero (hero):

you would multiply (3-x) by 6x/6x, yes correct

OpenStudy (anonymous):

you multiply the 6 by 7/7?

OpenStudy (anonymous):

wouldn't it give an improper fraction?

hero (hero):

You would get the following: \[\frac{7}{7} \times \frac{6}{1} = \frac{42}{7}\]

hero (hero):

Then you would add that to 6/7 to get: \[\frac{42}{7} + \frac{6}{7} = \frac{48}{7}\]

hero (hero):

But yes, 42/7 is an improper fraction. Notice that 42/7 = 6 But putting 6 in the proper fraction form with common denominator 7 is the only way to add it to fraction 6/7

OpenStudy (anonymous):

Got it. Can you help me with the this one: g/f(x)

hero (hero):

(g/f)(x) = g(x)/f(x)

hero (hero):

That's the general formula for dividing functions.

OpenStudy (anonymous):

1/x / 2-x?

hero (hero):

Yes, but to make it easier on yourself, write it like this: \[1/x \div 2 - x\]

hero (hero):

This way, you automatically know to inverse 2 - x

hero (hero):

\[\frac{1}{x} \times \frac{1}{2-x}\]

OpenStudy (anonymous):

Why is it multiplication now?

hero (hero):

Suppose we just had numbers:  Then what would you do?

OpenStudy (anonymous):

multiply the 3 by 2/2

hero (hero):

We're dividing not adding, so different rules apply. You should go back to learning about taking reciprocals of fractions when you're dividing by them.

OpenStudy (anonymous):

but would it be 1/2x-x^2